OpenStudy (anonymous):

In a certain country, there are 44.9 x 10^5 pet households, each owning 2.0 dogs on average. How many total dogs live in pet households?

OpenStudy (anonymous):

\[2\times44.9\times 10^5=89.8\times10^5=8980000\]\]